A bus left the rank carrying 64 passengers. At Taung ,k passengers got off the bus and 7 got in.
a.(i) Write an expression ,in terms of k for the number of passengers on the bus when it left Taung?
(ii) At Sebele, one third of the passengers got off the bus and 3 got on.
Write an expression in terms of k for the number of passengers on the bus when it left Sebele?
b.There are now 25 passengers on the bus.
Form an equation in terms of k,and solve it ?
c. Calculate the number of people on the bus after it left Taung?

Grade:10

1 Answers

Agrata Singh
5 years ago
a. K people got off and 7 got in
Number of people when bus leaves Taung=64-k+7=71-k
 
(ii).1/3 people got off. 1/3(71-k) got off. This means 2/3(71-k) people were left on bus. 3 more people got on
Total no. of people when bus leaves Sebele 2/3(71-k)+3 
 
b. Now 25 people on bus.
2/3(71-k)+3=25
2/3(71-k)=22
2(71-k)=22x3=66
(71-k)=66/2=33
K=71-33=38
Hence k = 38
 
c. When bus left Taung no of people=71-k=71-38=33
 
Hope this helps